An industry faces an inverse demand curve given by: P = 90 - 1.5Q. Let MC = 0.5Q before a new environmental-protection regulation and MC=10 + 0.5Q after the regulation.

Determine

a) The cost to consumers from the regulation

b) The social cost of the new regulation
Answer #1

a)

P = MC

90 - 1.5Q = 0.5Q

90 = 2Q

Q = 90/2

= 45

P =90 -1.5(45)

= 90 -67.5

= 22.5

Output and price after imposition of environmental regulation

90 - 1.5Q =10 + 0.5Q

80 = 2Q

Q = 40

P = 90 - 1.5(40)

= 90 -60

= 30

Loss to consumer:

Loss of Consumer Surplus = 0.5(Base)(Height)

= 0.5 (45)(90-22.5) - 0.5(40)(90-30)

= 318.75

b)

Loss to society:

= 0.5(5)(7.5)

=18.75

=18.75*2

=37.5